An ideal undergoes a reversible isothermal expansion at 132ºC. The entropy of the gas increase by 46.2 J/K. How much heat is absorbed?

Grade:upto college level

1 Answers

Deepak Patra
8 years ago
To obtain the absorbed heat Q, substitute 46.2 J/K for ΔS and 132̊ C for T in the equation Q = (ΔS) (T),
Q = (ΔS) (T)
= (46.2 J/K) (132̊ C)
= (46.2 J/K) ((132 +273) K)
= (46.2 J/K) (405 K)
= 1.87×104 J
From the above observation we conclude that, the absorbed heat Q will be 1.87×104 J.